OpenStudy (annie12m):

I'll medal and fan! If 65 1/8 was 100%, then what would be the percentage of 25.875? Show how you got the answer.

OpenStudy (study312):

well I would guess 0.100 x 25.875 = ? What is 0.100 x 25.875?

OpenStudy (annie12m):

2.5875

OpenStudy (study312):

Good now that in percent form is 258.75% that's by multiplying it by 100. Get it?

OpenStudy (annie12m):

idk i guess

OpenStudy (anonymous):

\[\frac{ 25.875 }{ 65\frac{ 1 }{ 8 } }\times 100=?\]

OpenStudy (annie12m):

25 7/8 = 25.875 and now ill make that into a mix number and same goes for 65 1/8 so 25 7/8 as a mix number = 207/8 and 65 1/8 as a mix number = 521/8 so it is 207/8 divided by 521/8 waitt but the second number is bigger than the first, so is it suppoe to be the other way around??

OpenStudy (annie12m):

@surjithayer

OpenStudy (anonymous):

\[\frac{ 207 }{ 8 }\times \frac{ 8 }{ 521 }\times 100=?\]